OpenStudy (anonymous):

solve for substitution ...PLEASE x-8y=-24 2x-2y=-6

OpenStudy (lgbasallote):

Hint: x = 24 + 8y

OpenStudy (lgbasallote):

x = -24 + 8y i mean

OpenStudy (anonymous):

x=24+8y 2(24+8y)-2y=-6 48+18y-2y=-6 48+16y=-6 16y=-54 y=-3.375